The sum of 3 numbers in an A.P is 21 and their product is 280. find the numbers

Answer by mananth(16946)

let the numbers be
a-d, a , a+d
sum = 21
a-d+a+a+d=21
3a=21
a=7
(a-d)(a)(a+d)=280
(a^2-d^2)*a=280
(49-d^2)*7=280
49-d^2= 280/7
49-d^2= 40
d^2=49-40
d^2=9
d= +/-3
(a-d),a,(a+d)
if d=3
4,7,10
if a=-3
10,7,4
(Academic)